OpenStudy (anonymous):

Find the first six terms of the sequence. (1 point) a1 = -8, an = 5 • an-1 -8, -40, -200, -1000, -5000, -25,000 -8, -40, -35, -30, -25, -20 0, 5, -40, -35, -30, -25 -40, -200, -1000, -5000, -25,000, -125,000 3. Find an equation for the nth term of the arithmetic sequence. (2 points) 8, 6, 4, 2, ... an = 8 + -2(n) an = 8 - 2 an = 8 + -2(n + 1) an = 8 + -2(n - 1) 4. Find an equation for the nth term of the arithmetic sequence. (2 points) -20, -16, -12, -8, ... an = -20 x 4(n - 1) an = -20 + 4(n - 1) an = -20 + 4(n + 2) an = -20 + 4(n + 1)

OpenStudy (welshfella):

for the first question each term is formed by multiplying the previous by 5

OpenStudy (welshfella):

- except for the first term a1 which is -8

OpenStudy (anonymous):

-8, -40, -35, -30, -25, -20 ??

OpenStudy (welshfella):

an means the nth term and an-1 means the term before it

OpenStudy (welshfella):

no does -40 * 5 = -35? you MULTIPLY by 5

OpenStudy (anonymous):

-8, -40, -200, -1000, -5000, -25,000 ??

OpenStudy (welshfella):

right 5 . an-1 means 4 times an-1

OpenStudy (welshfella):

* 5 times

OpenStudy (anonymous):

thank you ! could you also help me with 3 and 4?

OpenStudy (welshfella):

for an arithmetic series the nth term an:- an = a1 + (n-1)d where a1 = first term and d = common difference

OpenStudy (welshfella):

the common difference = second term - first term what is that in question 3?

OpenStudy (anonymous):

an = 8 + -2(n - 1) !

OpenStudy (welshfella):

yes a1 = 8 and d = -2

OpenStudy (welshfella):

now do number 4 is same way a1 = -20 d = -16 - -20 = ?

OpenStudy (welshfella):

Note the double negative

OpenStudy (anonymous):

-36

OpenStudy (welshfella):

no - - = +

OpenStudy (anonymous):

+4

OpenStudy (welshfella):

right

OpenStudy (welshfella):

so an = -20 + ?

OpenStudy (anonymous):

an = -20 + 4(n - 1) ??

OpenStudy (welshfella):

right!

OpenStudy (anonymous):

thank you so much :)

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

love you <3

OpenStudy (welshfella):

lol - dont tell my wife u said that!